OpenStudy (anonymous):

if AD=2x+3 and CD=x+10, what is the value of x?

OpenStudy (anonymous):

I think something's missing from this question. What do AD and CD represent?

OpenStudy (anonymous):

they are congruent

OpenStudy (anonymous):

answers (a)9 (b)11 (c)7 (d)15

OpenStudy (anonymous):

AD = CD 2x + 3 = AD x + 10 = CD Therefore 2x + 3 = x + 10 Solve for x.

OpenStudy (anonymous):

OK, so if AD=CD then sent the two statements equal \[2x+3=x+10\] and solve for x

OpenStudy (anonymous):

Move the 1x to the other side, and the 3 to the other side and the answer presents itself.

OpenStudy (anonymous):

is it 7?

OpenStudy (anonymous):

Very good.

OpenStudy (anonymous):

Yes